Exercise: Suppose the demand curve is linear and is given by the equation P = a – bQ where P is price and Q is quantity. What is the consumer surplus if the equilibrium price is P* and equilibrium quantity is Q*?


1
Expert's answer
2021-07-26T14:11:01-0400

Consumer Surplus = Total Utility – (Price x Quantity)


"Demand is given by : p = a - bQ"

"Marginal revenue(MR) = \frac{d(TR)}{dQ}"

"where TR = p*Q = (a - bQ)*Q"

"= aQ - bQ^2"

"Thus, MR = \frac{d(TR)}{dQ} = a - 2bQ"

"Hence, MR = a - 2bQ"

"When Q = 0 ,"

"then, P = a - b*0 = a"

"MR = a - 2b*0 = a"

Consumer surplus = "aQ\u2212bQ ^\n2 - (a \u00d7 0)"

"= aQ\u2212bQ ^\n2"

= 0-0 = 0